OpenStudy (anonymous):

simplify: (-5xy^2)-(-3xy^2 and ill pay you

OpenStudy (anonymous):

Don't the terms just add?

OpenStudy (anonymous):

combine like terms by computing \[-5-(-3)=-5+3=-2\]

OpenStudy (anonymous):

no charge

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

what kind of car is that?

OpenStudy (anonymous):

2 dollars

OpenStudy (anonymous):

73 Mustang

OpenStudy (anonymous):

nice. is it yours?

OpenStudy (anonymous):

Sure is.

OpenStudy (anonymous):

guess what i drive?

OpenStudy (anonymous):

73 satellite.

OpenStudy (anonymous):

not to hard to guess that one. also have a 66 belvedere, but i only drive it a couple times a month

OpenStudy (anonymous):

Nice.